Matlab Code if available Problem 04 mm us(0 In the circuit, the voltage source Vs(t) frequency, amplitude and phase are 22 kHz, 10 V and 0.25 rad corr
Answer #1

Vset solni- Cuven the electric circuit as 2 osoa 2 2 2 woon oa 05 = 138.23 Kre = 14,3240 V (H) ob Cuven into; - Volt) . . Vm=Rth = & Ron= Therudux voltage & vab = 2,486 1-88.93 so its amplitude – 2.4.86 volt. & praxe – – 880130 = 93.07° phase – 1.589V kth= (12.474) + 18.469 i) + (2-35.17) - (Rth=Rab] (12.474+18.46git 235.17) Rth = 120.4327 - 27,550) (14.474 +813.299) 8th =au registon 2 80 the penies & indectors connection - - R-3.560 X = 75,1772 fwl= 75, 177 r. - 5. 177 m . 2 38, 23 L - 000374 msry I have not available of matlab .